# Note on the Humbert function $\Psi_{1}$

### v1

@@
Abstract.
Recently, Hang and Luo (2024) used an elementary approach to derive the leading asymptotic behavior of the Humbert function $\Psi_{1}[a,b;c,c^{\prime};x,y]$ when $x\to\infty$ and $y\to+\infty$. In this paper, we give detailed asymptotic analyses of $\Psi_{1}$ when $x\to\infty$ and $y\to\infty$. Our proof is based on a uniform estimate of ${}_{2}F_{2}[a,b-n;c,d-n;z]$ for $z$ large.
Key words and phrases:
Humbert function, asymptotics, generalized hypergeometric function
2020 Mathematics Subject Classification:
33C20, 33C70, 41A60, 44A20
@@ 1. Introduction
The confluent hypergeometric function of two variables $\Psi_{1}$, introduced by Humbert [5, p. 75], is defined by
$\Psi_{1}[a,b;c,c^{\prime};x,y]=\sum_{m,n=0}^{\infty}\frac{(a)_{m+n}(b)_{m}}{(c)_{m}(c^{\prime})_{n}}\frac{x^{m}}{m!}\frac{y^{n}}{n!},\qquad|x|<1,\quad|y|<\infty,$
where $a,b\in\mathbb{C}$ and $c,c^{\prime}\notin\mathbb{Z}_{\leqslant 0}$. It can be analytically extended to the region
$\mathbb{D}_{\Psi_{1}}:=\bigl{\{}(x,y)\in\mathbb{C}^{2}\colon x\neq 1,\,|{\arg}(1-x)|<\pi,\,|y|<\infty\bigr{\}}$
via the series representation ([5, p. 75]; see also [3, equation (1.2)])
$\Psi_{1}[a,b;c,c^{\prime};x,y]=\sum_{n=0}^{\infty}\frac{\left(a\right)_{n}}{\left(c^{\prime}\right)_{n}}\,_{2}F_{1}\left[\begin{matrix}a+n,b\\ c\end{matrix};x\right]\frac{y^{n}}{n!},$
where ${}_{2}F_{1}$ denotes the Gauss hypergeometric function defined in (1.5).
By using inverse Laplace transformations of the Humbert functions $\Phi_{2}$, $\Phi_{3}$ and $\Xi_{2}$, Wald and Henkel [11] derived the leading-order behavior of these functions when two variables become simultaneously large. But they failed to render the integral for $\Psi_{1}$ [11, equation (2.4b)]
@@
as convolutions and so could not establish the asymptotics of $\Psi_{1}$.
Recently, Hang and Luo [3] partially answered Wald and Henkel’s problem by deriving the leading behavior of $\Psi_{1}$ under the condition
$x\to\infty,y\to+\infty,\quad\left|\arg(1-x)\right|<\pi,\quad 0<\gamma_{1}\leqslant\frac{y}{|1-x|}\leqslant\gamma_{2}<\infty.$ (1.1)
Their starting point is the following expression for $\Psi_{1}$.
Theorem 1.1 ([3, Theorem 3.1]).
@@
In this paper, we shall give detailed asymptotic analyses of $\Psi_{1}$ when $x\to\infty$ and $y\to\infty$. Our first result provides the full asymptotic expansion of $\Psi_{1}$ under the condition (1.1), which covers Hang and Luo’s result [3, Theorem 3.6].
@@
where $a_{0}(x,y)=1$ and in general,
$a_{k}(x,y)=\sum_{j=0}^{k}\frac{\left(b\right)_{j}\left(b-a+c^{\prime}\right)_{k-j}\left(j+b-a+1\right)_{k-j}}{j!(k-j)!}{}_{3}F_{2}\left[\begin{matrix}-j,j-k,c+c^{\prime}-a-1\\ b-a+c^{\prime},a-b-k\end{matrix};1\right]\Big{(}\frac{y}{1-x}\Big{)}^{b+j}.$ (1.3)
@@
$x\to\infty,y\to\infty,\quad\left|\arg(1-x)\right|<\pi,\quad\left|\arg(-y)\right|<\pi,\quad 0<\gamma_{1}\leqslant\Big{|}\frac{y}{1-x}\Big{|}\leqslant\gamma_{2}<\infty$ (1.4)
@@
$A_{2}(x,y)={}$ $\sum_{k=0}^{M}\frac{\left(c-b\right)_{k}}{k!}\left(-y\right)^{-a-k}\sum_{j=0}^{k-1}\frac{\left(b\right)_{j}\left(a-b\right)_{k-j}\left(a-b-c^{\prime}+1\right)_{k-j}}{j!\left(c-b\right)_{k-j}}\Big{(}\frac{y}{x-1}\Big{)}^{b+j},$
$A_{3}(x,y)={}$ $\sum_{k=0}^{M}\frac{\left(b\right)_{k}\left(c-b\right)_{k}}{k!^{2}}\left(-y\right)^{-a-k}{}_{2}F_{3}\left[\begin{matrix}b+k,b-c+1\\ 1+k,b-a+1,b-a+c^{\prime}\end{matrix};\frac{y}{1-x}\right]\Big{(}\frac{y}{x-1}\Big{)}^{b+k},$
@@
The paper are organised as follows. In Section 2, we demonstrate some estimates of the generalized hypergeometric functions ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ when $z$ approaches to infinity in any way, which are valid uniformly for $n\in\mathbb{Z}_{\geqslant 0}$. In Section 3, we use the tools in Section 2 to prove our main results. The paper concludes with further remarks in Section 4.
Notation. The number $C$ generically denotes a positive constant independent of the parameter $n$ and the variable $z$. By $f(n,z)=\mathcal{O}(a_{n}g(z))\,(z\in\Omega)$, we mean that there exists a constant $K>0$ independent of $n$ and $z$ such that
@@ 2. Auxiliary results
In this section, we prove some auxiliary results, including three lemmas about explicit bounds for the ratio of two gamma functions: (i) the first is benifitial to bound the tail series; (ii) the second generalizes the simple bound of gamma ratio [8, equation (5.6.8)]; (iii) the third guarantees the uniformity of explicit bounds for ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ obtained in the ensuing theorems.
Lemma 2.1 ([3, Lemma 2.1]).
@@
Proof.
Assume that $\Re(b-a)\geqslant 1,\,\Re(a)\geqslant 0$ and $\Re(z)>|\Im(a)|$. Then the quotient of two gamma functions has the integral representation [9, p. 33]
@@
Note that
@@
$=\frac{\Gamma(\Re(b-a))}{|\Gamma(b-a)|}\mathrm{e}^{\frac{\pi}{2}|\Im(a-b)|}\left(|z|+\Re(a)\cos\theta+\Im(a)\sin\theta\right)^{\Re(a-b)}.$
This completes the proof. ∎
@@
Proof.
Suppose that $n\geqslant 1$ and $z\in\bigcup_{j=1}^{n}\{|z+b-j|\geqslant\varepsilon\}$ for some $\varepsilon\in(0,\frac{1}{4})$. Denote
@@
$|-b+j-z|\geqslant|-b+j-z_{0}|.$
@@
$P_{n}(z)$ $\leqslant\prod_{j=1}^{n}\left(1+\frac{|a-b|}{|-b+j-z_{0}|}\right)=C\prod_{\begin{subarray}{c}1\leqslant j\leqslant n\\ |j-k|\geqslant 2\end{subarray}}\left(1+\frac{|a-b|}{\left|j-k-\varepsilon\,\mathrm{e}^{\mathrm{i}\beta}\right|}\right)$
@@
Assume that $a,b\in\mathbb{C},\,c\in\mathbb{C}\setminus\mathbb{Z}_{\geqslant 0}$ and $d,a-b\in\mathbb{C}\setminus\mathbb{Z}$. Let $w>0$ be a number such that $w>\max\{\Re(a),\Re(b),\Re(d)\}$ and that the fractional parts of $w-\Re(a)$ and $w-\Re(b)$ are both in the interval $(\varepsilon,1)$, where $\varepsilon>0$ is a small number. Then for fixed $n\in\mathbb{Z}_{\geqslant 0}$,
@@
and the contours $L^{\pm}$ which pass to infinity in the directions ±$\pm\theta_{0}\,(0<\theta_{0}<\frac{\pi}{2})$. Further, $\mathcal{C}$ is taken to embrace all the poles of $\Gamma(-s)$ and the points
@@
The convergence of the integral as $|s|\to\infty$ is controlled by the term $\exp\{-|s|\cos\theta_{0}\cdot\log|s|\}$ in the behaviour of the gamma function for large $|s|$. Use of Cauchy’s residue theorem implies that
@@
See [6, Section 2] for details. Divide $R_{n,w}(z)$ into two parts:
@@
$\left|\frac{\Gamma(a-w+\mathrm{i}t)}{\Gamma(c-w+\mathrm{i}t)}\right|=\mathcal{O}(1),\quad\left|\Gamma(-w-\mathrm{i}t)\right|=\mathcal{O}(1).$ (2.5)
@@
If $|t|\leqslant\max\{1,|\Im(d)|\}$, apply [8, equations (5.6.6) and (5.6.7)] and (2.1) to yield
@@
If $|t|>\max\{1,|\Im(d)|\}$, take $r>0$ such that $w>r>\Re(d)$. Put $z_{1}=(1+w+n-r)-\mathrm{i}(t+\Im(d))$ and then
@@
When $\Re(b_{1}-a_{1})\geqslant 1$, Lemma 2.2 shows that
@@
When $\Re(b_{1}-a_{1})<1$, put $A=\lceil\Re(a_{1}-b_{1})\rceil$ and $a_{1}^{\prime}=a_{1}-A-1$. Then $\Re(b_{1}-a_{1}^{\prime})\geqslant 1$ and further,
@@
Step 2. Estimate of $R_{\pm}(z)$. Due to [9, Lemma 2.1] and Euler’s reflection formula, we obtain that for $s\in L^{\pm}$,
$h_{0}(s)=\frac{\Gamma(a+s)\Gamma(b+s)}{\Gamma(c+d)\Gamma(d+s)\Gamma(1+s)}\frac{-\pi}{\sin(\pi s)}=\frac{-\pi\csc(\pi s)\rho_{0}(s)}{\Gamma(s+1-\alpha)}=\rho_{0}(s)\Gamma(-s+\alpha)\frac{\sin(\pi(s-\alpha))}{\sin(\pi s)},$
@@
$\left|\frac{\sin(\pi(s-\alpha))}{\sin(\pi s)}\right|=\mathcal{O}(1),\quad s\in L^{\pm}.$
Therefore, for $s\in L^{\pm}$, there are positive constants $K_{1}$ and $K_{2}$ independent of $n$ and $z$, such that when $|s|\geqslant K_{1}$,
$|\rho_{0}(s)|\leqslant K_{2},\quad\left|\frac{\sin(\pi(s-\alpha))}{\sin(\pi s)}\right|\leqslant K_{2}.$
@@
$\left|R_{\pm}(z)\right|$ $\leqslant C\left(n+1\right)^{2|b-d|}\left(\int_{s\in L^{\pm},|s|<K_{1}}|h_{0}(s)||ds|+\int_{s\in L^{\pm},|s|\geqslant K_{1}}|h_{0}(s)||ds|\right)$
@@
According to [9, Lemma 2.8], the contours $L^{\pm}$ can be deformed to pass through the saddle point $s=-z$ and to be bounded away from the points shown in (2.4). Hence, the integral in the last line has the order of magnitude $\mathcal{O}(|z|^{\alpha}\mathrm{e}^{-\Re(z)})$.
The expansion (2.2) follows from the estimates of $R_{v}(z)$ and $R_{\pm}(z)$. ∎
@@
as $z\to\infty$ such that $\left|\arg(z)\right|<\pi$ and $z$ is bounded way from the points $-b+k\,(k\in\mathbb{Z})$.
Proof.
@@
where $R_{-n,0}(z)$ is defined in (2.3). The result follows by repeating the proof of Theorem 2.4. ∎
The third establishes explicit upper bounds of ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ for $z\to+\infty$ and refines the rough bounds in [3, Lemma 3.3].
@@
Proof.
By follwing the proof of [3, Lemma 3.3], the results follow from Lemma 2.3. ∎

### v3

# Complete asymptotic expansions of the Humbert function $\Psi_{1}$ for two large arguments

@@
Abstract
In our recent work [SIGMA 20 (2024), 074, 13 pages], the leading behaviour of the Humbert function $\Psi_{1}[a,b;c,c^{\prime};x,y]$ when $x\to\infty$ and $y\to+\infty$ has been derived in a direct and simple manner. In this paper, we obtain the complete asymptotics of $\Psi_{1}$ in the general case $x,y\to\infty$ along a new path. Indeed, our proof is based on a sharp estimate on ${}_{2}F_{2}[a,b-n;c,d-n;z]$, which is valid uniformly for $n\in\mathbb{Z}_{\geqslant 0}$ and large $z$.
MSC 2020: primary 33C70; 41A60; secondary 30E15
Keywords: Humbert functions; Asymptotic expansions
@@ 1.  Introduction
The confluent hypergeometric function of two variables $\Psi_{1}$, introduced by Humbert [6, p. 75], is defined by
$\Psi_{1}[a,b;c,c^{\prime};x,y]:=\sum_{m,n=0}^{\infty}\frac{(a)_{m+n}(b)_{m}}{(c)_{m}(c^{\prime})_{n}}\frac{x^{m}}{m!}\frac{y^{n}}{n!},\qquad|x|<1,|y|<\infty,$
where $a,b\in\mathbb{C}$ and $c,c^{\prime}\notin\mathbb{Z}_{\leqslant 0}$. It is known that $\Psi_{1}$ has an extension to the region (see [4, 5])
$\mathbb{D}_{\Psi_{1}}:=\bigl{\{}(x,y)\in\mathbb{C}^{2}\colon x\neq 1,\,\left|\arg(1-x)\right|<\pi,\,|y|<\infty\bigr{\}}.$ (1.1)
By using inverse Laplace transformations of the Humbert functions $\Phi_{2}$, $\Phi_{3}$ and $\Xi_{2}$, Wald and Henkel [12] derived the leading-order behaviour of these functions when two variables become simultaneously large. But they failed to interpret the integral for $\Psi_{1}$ [12, Equation (2.4b)]
@@
as a convolution and so could not establish the asymptotics of $\Psi_{1}$.
Recently, we [4] partially answered Wald and Henkel’s problem by deriving the leading behaviour of $\Psi_{1}$ under the condition
$x\to\infty,y\to+\infty,\quad\left|\arg(1-x)\right|<\pi,\quad 0<\gamma_{1}\leqslant\frac{y}{\left|1-x\right|}\leqslant\gamma_{2}<\infty.$ (1.2)
Our starting point is the following expression for $\Psi_{1}$.
Theorem 1.1 ([4, Theorem 3.1]).
@@
In this paper, we shall give detailed asymptotic analyses of $\Psi_{1}$ when $x\to\infty$ and $y\to\infty$. Our first result provides the full asymptotic expansion of $\Psi_{1}$ under the condition (1.2), which covers our previous result [4, Theorem 3.6].
@@
where $a_{0}(x,y)=\bigl{(}\frac{y}{1-x}\bigr{)}^{b}$ and in general, for any $k\in\mathbb{Z}_{\geqslant 0}$,
$a_{k}(x,y)=\sum_{j=0}^{k}\frac{\left(b\right)_{j}\left(b-a+c^{\prime}\right)_{k-j}\left(j+b-a+1\right)_{k-j}}{j!\left(k-j\right)!}{}_{3}F_{2}\left[\begin{matrix}-j,j-k,c+c^{\prime}-a-1\\ b-a+c^{\prime},a-b-k\end{matrix};1\right]\left(\frac{y}{1-x}\right)^{b+j}.$ (1.4)
@@
$x\to\infty,y\to\infty,\quad\left|\arg(1-x)\right|<\pi,\quad\left|\arg(-y)\right|<\pi,\quad 0<\gamma_{1}\leqslant\left|\frac{y}{1-x}\right|\leqslant\gamma_{2}<\infty$ (1.5)
@@
$A_{2}(x,y)={}$ $\sum_{k=0}^{M}\frac{\left(c-b\right)_{k}}{k!}\left(-y\right)^{-a-k}\sum_{j=0}^{k-1}\frac{\left(b\right)_{j}\left(a-b\right)_{k-j}\left(a-b-c^{\prime}+1\right)_{k-j}}{j!\left(c-b\right)_{k-j}}\left(\frac{y}{x-1}\right)^{b+j},$
$A_{3}(x,y)={}$ $\sum_{k=0}^{M}\frac{\left(b\right)_{k}\left(c-b\right)_{k}}{\left.k!\right.^{2}}\left(-y\right)^{-a-k}{}_{3}F_{3}\left[\begin{matrix}b+k,b-c+1,1\\ 1+k,b-a+1,b-a+c^{\prime}\end{matrix};\frac{y}{1-x}\right]\left(\frac{y}{x-1}\right)^{b+k},$
@@
The paper is organized as follows. In Section 2, we demonstrate some estimates of the generalized hypergeometric functions ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ when $z$ goes all the way to infinity, which are valid uniformly for $n\in\mathbb{Z}_{\geqslant 0}$. In Section 3, we use the tools in Section 2 to prove our main results. The paper concludes with further remarks in Section 4.
Notation. The number $C$ generically denotes a positive constant independent of the summation index $n$ and the variable $z$. By $f(n,z)=\mathcal{O}(a_{n}g(z))\,(z\in\Omega)$, we mean that there exists a constant $K>0$ independent of $n$ and $z$ such that
@@ 2.  Auxiliary results
In this section, we prove some auxiliary results, including three lemmas about explicit bounds for the ratio of two gamma functions: (i) the first is beneficial to bound the tail series; (ii) the second generalizes the simple bound of gamma ratio [9, Equation (5.6.8)]; (iii) the third guarantees the uniformity of explicit bounds for ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ obtained in the ensuing theorems.
Lemma 2.1 ([4, Lemma 2.1]).
@@
Proof..
Recall the integral representation for the quotient of two gamma functions [10, p. 33]: for $b>a\geqslant 0$ and $\Re(z)>0$,
@@
By analytic continuation, it is valid for $\Re(b)>\Re(a)\geqslant 0$ and $\Re(z)>|\Im(a)|$. Note that
@@
$=\frac{\Gamma(\Re(b-a))}{|\Gamma(b-a)|}\mathrm{e}^{\frac{\pi}{2}|\Im(a-b)|}\left(|z|+\Re(a)\cos\theta+\Im(a)\sin\theta\right)^{\Re(a-b)},$
where the identity
$\int_{0}^{\infty}u^{\alpha-1}\mathrm{e}^{-\lambda u}\mathrm{d}u=\frac{\Gamma(\alpha)}{\lambda^{\alpha}}~{}~{}~{}(\min\{\Re(\alpha),\Re(\lambda)\}>0)$
is used. ∎
@@
Proof..
Suppose that $n\geqslant 1$ and $z\in\bigcup_{j=1}^{n}\big{\{}|z+b-j|\geqslant\varepsilon\big{\}}$ for some $\varepsilon\in\big{(}0,\frac{1}{4}\big{)}$. Denote
@@
$\left|-b+j-z\right|\geqslant\left|-b+j-z_{0}\right|.$
@@
$P_{n}(z)$ $\leqslant\prod_{j=1}^{n}\left(1+\frac{|a-b|}{\left|-b+j-z_{0}\right|}\right)=C\prod_{\begin{subarray}{c}1\leqslant j\leqslant n\\ |j-k|\geqslant 2\end{subarray}}\left(1+\frac{|a-b|}{\left|j-k-\varepsilon\,\mathrm{e}^{\mathrm{i}\beta}\right|}\right)$
@@
Assume that $a,b\in\mathbb{C},\,c\in\mathbb{C}\setminus\mathbb{Z}_{\geqslant 0}$ and $d,a-b\in\mathbb{C}\setminus\mathbb{Z}$. Let $w>0$ be a number such that $w>\max\{\Re(a),\Re(b),\Re(d)\}$ and that the fractional parts of $w-\Re(a)$ and $w-\Re(b)$ are both in the interval $(\varepsilon,1)$, where $\varepsilon>0$ is a small number. Then for any $n\in\mathbb{Z}_{\geqslant 0}$,
@@
and the contours $L^{\pm}$ which pass to infinity in the directions ±$\pm\theta_{0}$ $\big{(}0<\theta_{0}<\frac{\pi}{2}\big{)}$. Further, $\mathcal{C}$ is taken to embrace all the poles of $\Gamma(-s)$ and the points
@@
The convergence of the integral as $|s|\to\infty$ is guaranteed by the term $\exp\{-|s|\cos\theta_{0}\cdot\log|s|\}$ in the behaviour of the gamma function for large $|s|$. The use of Cauchy’s residue theorem implies that
@@
See [7, Section 2] for details. Divide $R_{n,w}(z)$ into two parts:
@@
$\left|\frac{\Gamma(a-w+\mathrm{i}t)}{\Gamma(c-w+\mathrm{i}t)}\Gamma(-w-\mathrm{i}t)\right|=\mathcal{O}(1).$ (2.5)
@@
If $|t|\leqslant\max\{1,|\Im(d)|\}$, apply [9, Equations (5.6.6) and (5.6.7)] and (2.1) to yield
@@
If $|t|>\max\{1,|\Im(d)|\}$, take $r>0$ such that $w>r>\Re(d)$. Set $z_{1}=(1+w+n-r)-\mathrm{i}(t+\Im(d))$ and then
@@
When $\Re(b_{1}-a_{1})=\Re(d-b)\geqslant 1$, Lemma 2.2 shows that
@@
When $\Re(b_{1}-a_{1})<1$, let $A=\lceil\Re(a_{1}-b_{1})\rceil$ and $a_{1}^{\prime}=a_{1}-A-1$. Then $\Re(b_{1}-a_{1}^{\prime})\geqslant 1$ and further,
@@
Step 2. Estimate of $R_{\pm}(z)$. Due to [10, Lemma 2.1] and Euler’s reflection formula, we obtain that for $s\in L^{\pm}$,
$h_{0}(s)$ $=\frac{\Gamma(1-c-s)\Gamma(1-d-s)}{\Gamma(1-a-s)\Gamma(1-b-s)}\Gamma(-s)\cdot\frac{\sin(\pi(c+s))\sin(\pi(d+s))}{\sin(\pi(a+s))\sin(\pi(b+s))}$
$=\rho_{0}(-s)\Gamma(-s+\alpha)\cdot\frac{\sin(\pi(c+s))\sin(\pi(d+s))}{\sin(\pi(a+s))\sin(\pi(b+s))},$
@@
$\left|\frac{\sin(\pi(c+s))\sin(\pi(d+s))}{\sin(\pi(a+s))\sin(\pi(b+s))}\right|=\mathcal{O}(1),\quad s\in L^{\pm}.$
Therefore, for $s\in L^{\pm}$, there are positive constants $K_{1}$ and $K_{2}$ independent of $n$ and $z$, such that when $|s|\geqslant K_{1}$ then
$|\rho_{0}(-s)|\leqslant K_{2},\quad\left|\frac{\sin(\pi(c+s))\sin(\pi(d+s))}{\sin(\pi(a+s))\sin(\pi(b+s))}\right|\leqslant K_{2}.$
@@
$\left|R_{\pm}(z)\right|$ $\leqslant C\left(n+1\right)^{2|b-d|}\left(\int_{s\in L^{\pm},|s|<K_{1}}|h_{0}(s)z^{s}||ds|+\int_{s\in L^{\pm},|s|\geqslant K_{1}}|h_{0}(s)z^{s}||ds|\right)$
@@
According to [10, Lemma 2.8], the contours $L^{\pm}$ can be deformed to pass through the saddle point $s=-z$ and to be bounded away from the points shown in (2.4). Hence, the integral in the last line has the order of magnitude $\mathcal{O}\big{(}|z|^{\Re(\alpha)}\mathrm{e}^{-\Re(z)}\big{)}$.
The expansion now (2.2) follows from the estimates of $R_{v}(z)$ and $R_{\pm}(z)$. ∎
@@
as $z\to\infty$ such that $\left|\arg(z)\right|<\pi$ and $z$ is bounded away from the points $-b+k\,(k\in\mathbb{Z})$.
Proof..
@@
where $R_{-n,0}(z)$ is defined in (2.3) (The contour is suitably indented in order to avoid the pole $s=0$ of $\Gamma(-s)$). The result follows by repeating the proof of Theorem 2.4. ∎
The third establishes explicit upper bounds of ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ for $z\to+\infty$ and refines the rough bounds in [4, Lemma 3.3].
@@
Proof..
By following the proof of [4, Lemma 3.3], the results follow from Lemma 2.3. ∎